# The Hierarchical Growth of Bright Central Galaxies and Intracluster Light as Traced by the Magnitude Gap

### v1

@@ Abstract
Using a sample of 2800 galaxy clusters identified in the Dark Energy Survey across the redshift range , we characterize the hierarchical assembly of Bright Central Galaxies (BCGs) and the surrounding intracluster light (ICL). To quantify hierarchical formation we use the stellar mass - halo mass (SMHM) relation for the BCG+ICL system and incorporate the magnitude gap (M14), the difference in brightness between the BCG (measured within 30 kpc) and 4th brightest cluster member galaxy within 0.5 . The inclusion of M14, which traces BCG hierarchical growth, increases the slope and decreases the intrinsic scatter in the SMHM relation, highlighting that it is a latent variable within the BCG+ICL SMHM relation. Moreover, the correlation with M14 decreases at large radii from the BCG’s centre. However, the stellar light within the BCG+ICL transition region (30 kpc - 80 kpc) most strongly correlates with the dark matter halo mass and has a statistically significant correlation with M14. As the light in the transition region and M14 are independent measurements, the transition region may grow as a result of the BCG’s hierarchical two-phase formation. Additionally, as M14 and ICL result from hierarchical growth, we use a stacked sample and find that clusters with large M14 values are characterized by larger ICL and BCG+ICL fractions, which illustrates that the merger processes that build the BCG stellar mass also grow the ICL. Furthermore, this may suggest that M14 combined with the ICL fraction can be used as a method to identify dynamically relaxed clusters.
1 Introduction
@@
Bright Central Galaxies (BCGs) are massive, radially extended, elliptical galaxies located near the centre of their host cluster’s dark matter halo. BCGs grow hierarchically, by merging with smaller galaxies over time (e.g., De Lucia & Blaizot, 2007; Oser et al., 2010; van Dokkum et al., 2010). The combination of hierarchical formation and central location results in a correlation between the properties of the BCG and the cluster’s dark matter halo (Jones & Forman, 1984; Rhee & Latour, 1991; Lin & Mohr, 2004; Lauer et al., 2014). This galaxy-halo connection provides key information in characterizing the formation history of these massive galaxies.
@@
Observations, simulations, and semi-analytic models (Croton et al., 2006; De Lucia & Blaizot, 2007; Guo et al., 2011; Tonini et al., 2012; Shankar et al., 2015) suggest that BCGs grow as a result of a two-phase formation (Oser et al., 2010; van Dokkum et al., 2010). At high redshift () a dense core (10kpc), which contains 25% of the stellar mass is formed via in-situ star formation. At , the outer envelope grows hierarchically as a result of major and minor mergers. This outer envelope extends to the faint and diffuse halo of intracluster light (ICL; Zwicky, 1933, 1951) that is observed around BCGs. Indeed, the presence of ICL can be used to differentiate BCGs from similarly massive non-central galaxies. Therefore, it is unsurprising that prior works have suggested that at least part of the ICL forms due to the BCG’s hierarchical assembly process (e.g., Murante et al., 2007; Golden-Marx et al., 2023).
@@
Due to the faint and diffuse nature of the ICL, deep, high resolution photometry with a large field-of-view is needed to accurately characterize the ICL. Using such observations, individual clusters have measurements of the ICL that extend out to hundreds of kiloparsecs from the BCG (e.g., Kluge et al., 2020; Montes et al., 2021; Montes & Trujillo, 2022; Golden-Marx et al., 2023; Kluge et al., 2024). However, these observations can be enhanced by stacking the ICL signal surrounding many BCGs. Doing so, prior analyses have found that the ICL extends as far out as Mpc scales (Zibetti et al., 2005; Zhang et al., 2019c; Chen et al., 2022; Zhang et al., 2024). Although the ICL can be measured out to large radii, it is observationally challenging to distinguish the light associated with the BCG and that of the ICL. Therefore, we do not disentangle the BCG and ICL; instead, we focus on the BCG+ICL system and use fixed radial apertures to define our measurements, as introduced in Pillepich et al. (2018). In this analysis, we define similar radial regimes as Zhang et al. (2024); the BCG refers to the inner 30 kpc, the BCG+ICL transition regime is 30 kpc - 80 kpc, and the ICL is the light beyond 80 kpc.
@@
As a result of the two-phase growth, information about the BCG’s recent stellar mass growth is contained within the BCG’s outer envelope and ICL (Oser et al., 2010; van Dokkum et al., 2010). It follows that measurements of the SMHM relation in both observations and semi-analytic models have found that the slope increases in progressively larger radii (Moster et al., 2018; Golden-Marx & Miller, 2019). Moreover, Golden-Marx et al. (2023) extended this analysis to the ICL and found that the SMHM relation’s slope increases to an asymptote at 100 kpc. Moreover, recent work found that the underlying halo mass is more strongly correlated with the light from the outer envelopes (50 kpc-100 kpc (Huang et al., 2022) or 30 kpc - 50 kpc (Golden-Marx et al., 2023)) than the light from the BCG’s core. Thus, the ICL is imprinted with information about recent mergers, which strengthens the correlation with the underlying halo mass.
Accounting for latent variables, which ideally increase the slope and decrease also strengthens statistical correlations. One observational measurement inherently tied to BCG hierarchical growth is the magnitude gap, the difference in band magnitude between the BCG and 4th brightest cluster member within half the radius enclosing 200 times the critical density of the Universe () (Dariush et al., 2010), referred to as M14 throughout this analysis. N-body simulations found that BCG stellar mass linearly increases with the number of progenitor galaxies (Solanes et al., 2016). Thus, as BCGs grow hierarchically, they increase in stellar mass and brightness, while the satellite galaxies remain fixed or merge with the BCG. Therefore, BCG growth increases M14 yielding the correlation between M14 and BCG stellar mass (Harrison et al., 2012; Golden-Marx & Miller, 2018, 2019; Golden-Marx et al., 2022). Thus, M14 traces the BCG’s hierarchical assembly (Golden-Marx & Miller, 2018) and acts as an indicator of the BCG’s dominance over other cluster member galaxies, similar to the Bautz-Morgan classification (Bautz & Morgan, 1970). Additionally, Vitorelli et al. (2018) found that clusters with larger magnitude gaps, defined using the 2nd brightest galaxy (M12), are more concentrated and likely formed earlier, suggesting that the magnitude gap traces cluster formation.
@@ 2.1 DES Y6 data
The data used throughout this analysis, both the source catalogs and the images, come from the DES Y6 co-addition (this includes publicly released data from DES Data Release 2 (Abbott et al., 2021)). The DESY6 images are deeper than the Y3 versions and provide a more uniform co-add with an improved background subtraction. Since this work focuses on the diffuse ICL, these improvements, which include the improved masking of faint galaxies located spatially near/within the ICL, are key. Such faint galaxies can unintentionally be treated as part of the ICL, thus, these improvements enhance our ability to accurately measure the contribution from the ICL by strengthening our ability to delineate between the diffuse light and total light within the cluster, which includes all cluster member galaxies and the ICL.
@@ 2.3 Measuring the ICL
We analyse only clusters in the redshift range 0.2 0.6, which reduces the number of redMaPPer ( 20) clusters from 21092 to 15654.
@@
Using these BCGs, the DES database was queried to select images within a 0.15 deg 0.15 deg region centred on the BCG. These individual images were combined (using mean pixel values) to create the final coadd image centred on the BCG. The images used throughout this analysis have not had a local sky background subtraction (measured either with Source Extractor (Bertin & Arnouts, 1996) or Swarp (Bertin, 2010)) applied. Instead, these images underwent a global background subtraction process (estimated across the whole Field-of-View, about 2.5 for single exposure images; Bernstein et al., 2017). This background subtraction is done because local background regions often include the ICL, which results in an underestimation of the light in the BCG+ICL outer profiles (e.g., von der Linden et al., 2007; Golden-Marx & Miller, 2018).
To prevent contamination of our BCG+ICL light profiles from bright foreground stars and galaxies, clusters with BCGs near such objects were removed, reducing the number of available clusters in our sample to 7042.
Using the BCG as the centre, all objects covering a 0.2 deg 0.2 deg region surrounding the BCG in the DES database were identified. We acknowledge that miscentring exists in redMaPPer (see Zhang et al. (2019a)). However, this is not corrected for in the reduction process. As shown in the Appendix, when we look at high and low probability centrals, we find no evidence that it impacts the parameters of the SMHM relation as shown in Tables 4 and 5.
Elliptical masks, with a semi-major axis of 3.5 (Kron, 1980), were placed around all galaxies above a selected masking limit, excluding the BCG, allowing for the measurement of a BCG+ICL profile. The masking limit is a -band magnitude limit, based on the cluster’s photometric redshift, of 0.2, where is the characteristic luminosity of a cluster red galaxy luminosity function measurement (Zhang et al., 2019b).
@@
Lastly, we remind the reader that unlike Zhang et al. (2019c), Sampaio-Santos et al. (2021), and Zhang et al. (2024), we do not stack the measurements of the ICL on a secondary cluster property such as richness. Instead, we only combine images of the same cluster to create a deeper photometric image.
@@ 3 The Observable Measurements for Constructing the SMHM Relation
3.1 BCG Stellar Mass
@@
Using the selected aperture magnitudes, we use the EzGal (Mancone & Gonzalez, 2012) SED modelling software to estimate the stellar mass. For EzGal, we assume a passive spectral model, as DES photometry prevents us from statistically constraining additional star formation parameters such as burst times and formation epochs. For converting magnitude to stellar mass, we assume the same parameters as Golden-Marx & Miller (2019), Golden-Marx et al. (2022), and Golden-Marx et al. (2023); a Bruzual & Charlot (2003) stellar population synthesis model, a Salpeter (1955) Initial Mass Funtion (IMF), a formation redshift of , and a metallicity of 0.008 (66% Z⊙). These parameters were selected to minimize the statistic between the measured and EzGal modeled photometry. Following Golden-Marx et al. (2022) and Golden-Marx et al. (2023), we assume a subsolar metallicity for each BCG+ICL system since observations find that the metallicity of early type galaxies decreases radially outward (McDermid et al., 2015) and light from the ICL has been characterized as having a subsolar metallicity (Montes & Trujillo, 2018). Additionally, we note that our stellar mass estimate is independent of our choice of formation redshift and that changing the Salpeter (1955) IMF to a Chabrier (2003) IMF uniformly lowers our stellar mass estimates by 0.25 dex. Thus, the SMHM relation’s amplitude is the only SMHM parameter impacted.
@@ 3.3 Magnitude Gap (M14)
M14 is defined uniformly throughout this paper, regardless of our choice of BCG+ICL aperture stellar mass. For the DES-redMaPPer clusters we identify cluster members as those with . This is a less restrictive than Golden-Marx & Miller (2019) and Golden-Marx et al. (2022) since we do not match to the C4 catalog (Miller et al., 2005). This choice optimizes the number of clusters in our sample and only reduces the sample from 7042 to 6903 clusters. When we did this analysis with a threshold of our results were consistent.

### v2

@@ Abstract
Using a sample of 2800 galaxy clusters identified in the Dark Energy Survey across the redshift range , we characterize the hierarchical assembly of Bright Central Galaxies (BCGs) and the surrounding intracluster light (ICL). To quantify hierarchical formation we use the stellar mass - halo mass (SMHM) relation, comparing the halo mass, estimated via the mass-richness relation, to the stellar mass within the BCG+ICL system. Moreover, we incorporate the magnitude gap (M14), the difference in brightness between the BCG (measured within 30 kpc) and 4th brightest cluster member galaxy within 0.5 , as a third parameter in this linear relation. The inclusion of M14, which traces BCG hierarchical growth, increases the slope and decreases the intrinsic scatter, highlighting that it is a latent variable within the BCG+ICL SMHM relation. Moreover, the correlation with M14 decreases at large radii. However, the stellar light within the BCG+ICL transition region (30 kpc - 80 kpc) most strongly correlates with halo mass and has a statistically significant correlation with M14. Since the transition region and M14 are independent measurements, the transition region may grow due to the BCG’s hierarchical formation. Additionally, as M14 and ICL result from hierarchical growth, we use a stacked sample and find that clusters with large M14 values are characterized by larger ICL and BCG+ICL fractions, which illustrates that the merger processes that build the BCG stellar mass also grow the ICL. Furthermore, this may suggest that M14 combined with the ICL fraction can identify dynamically relaxed clusters.
keywords:
galaxies: clusters: general – galaxies: elliptical and lenticular, cD – galaxies: evolution1 Introduction
@@
Bright Central Galaxies (BCGs) are massive, radially extended, elliptical galaxies located near the centre of their host cluster’s dark matter halo. BCGs grow hierarchically, by merging with smaller galaxies over time (e.g., De Lucia & Blaizot, 2007; Bezanson et al., 2009; Naab et al., 2009; Oser et al., 2010; van Dokkum et al., 2010; Wellons et al., 2016). The combination of hierarchical formation and central location results in a correlation between the properties of the BCG and the cluster’s dark matter halo (Jones & Forman, 1984; Rhee & Latour, 1991; Lin & Mohr, 2004; Lauer et al., 2014). This galaxy-halo connection provides key information in characterizing the formation history of these massive galaxies.
@@
Observations, simulations, and semi-analytic models (Croton et al., 2006; De Lucia & Blaizot, 2007; Guo et al., 2011; Tonini et al., 2012; Shankar et al., 2015) suggest that BCGs grow as a result of a two-phase formation (Oser et al., 2010; van Dokkum et al., 2010). At high redshift () a dense core (10kpc), which contains 25% of the stellar mass is formed via in-situ star formation. At , the outer envelope grows hierarchically as a result of major and minor mergers. This outer envelope extends to the faint and diffuse halo of intracluster light (ICL; Zwicky, 1933, 1951) that is observed around BCGs. Indeed, the presence of ICL can be used to differentiate BCGs from similarly massive non-central galaxies (e.g., Hoessel et al., 1980; Lynam et al., 2000). Therefore, it is unsurprising that prior works have suggested that at least part of the ICL forms due to the BCG’s hierarchical assembly process (e.g., Murante et al., 2007; Golden-Marx et al., 2023).
@@
Due to the faint and diffuse nature of the ICL, deep, high resolution photometry with a large field-of-view is needed to accurately characterize the ICL. Using such observations, individual clusters have measurements of the ICL that extend out to hundreds of kiloparsecs from the BCG (e.g., Kluge et al., 2020; Gonzalez et al., 2021; Montes et al., 2021; Montes & Trujillo, 2022; Golden-Marx et al., 2023; Kluge et al., 2024). However, these observations can be enhanced by stacking the ICL signal surrounding many BCGs. Doing so, prior analyses have found that the ICL extends as far out as Mpc scales (Zibetti et al., 2005; Zhang et al., 2019c; Chen et al., 2022; Zhang et al., 2024). Although the ICL can be measured out to large radii, it is observationally challenging to distinguish the light associated with the BCG and that of the ICL (see discussions in e.g., Gonzalez et al., 2007; Zhang et al., 2019c; Gonzalez et al., 2021; Kluge et al., 2021; Montes et al., 2021; Brough et al., 2024). Therefore, we do not disentangle the BCG and ICL; instead, we focus on the BCG+ICL system and use fixed radial apertures to define our measurements, as introduced in Pillepich et al. (2018). In this analysis, we define similar radial regimes as Zhang et al. (2024); the BCG refers to the inner 30 kpc, the BCG+ICL transition regime is 30 kpc - 80 kpc, and the ICL is the light beyond 80 kpc.
@@
As a result of the two-phase growth, information about the BCG’s recent stellar mass growth is contained within the BCG’s outer envelope and ICL (Oser et al., 2010; van Dokkum et al., 2010). It follows that measurements of the SMHM relation in both observations and semi-analytic models have found that the slope increases in progressively larger radii (Moster et al., 2018; Golden-Marx & Miller, 2019). Moreover, Golden-Marx et al. (2023) extended this analysis to the ICL and found that the SMHM relation’s slope increases to an asymptote at 100 kpc. Moreover, recent work found that the underlying halo mass is more strongly correlated with the light from the outer envelopes (50 kpc-100 kpc (Huang et al., 2022; Kwiecien et al., 2024) or 30 kpc - 50 kpc (Golden-Marx et al., 2023)) than the light from the BCG’s core. Thus, the ICL is imprinted with information about recent mergers, which strengthens the correlation with the underlying dark matter halo mass (e.g., Montes & Trujillo, 2019).
Accounting for latent variables, which ideally increase the slope and decrease also strengthens statistical correlations. One observational measurement inherently tied to BCG hierarchical growth is the magnitude gap, the difference in band magnitude between the BCG and 4th brightest cluster member within half the radius enclosing 200 times the critical density of the Universe () (Dariush et al., 2010), referred to as M14 throughout this analysis. N-body simulations found that BCG stellar mass linearly increases with the number of progenitor galaxies (Solanes et al., 2016). Thus, as BCGs grow hierarchically, they increase in stellar mass and brightness, while the satellite galaxies remain fixed or merge with the BCG. Therefore, BCG growth increases M14 yielding the correlation between M14 and BCG stellar mass (Harrison et al., 2012; Golden-Marx & Miller, 2018, 2019; Golden-Marx et al., 2022). Thus, M14 traces the BCG’s hierarchical assembly (Golden-Marx & Miller, 2018) and acts as an indicator of the BCG’s dominance over other cluster member galaxies, similar to the Bautz-Morgan classification (Bautz & Morgan, 1970). We also note that Oliva-Altamirano et al. (2014) identify a similar offset in stellar mass at fixed halo mass between central and non-central BCGs, which they find have low magnitude gaps, defined using the 2nd brightest galaxy (M12), in galaxy groups and clusters from the GAMA survey. Additionally, Vitorelli et al. (2018) found that clusters with larger M12 values, are more concentrated and likely formed earlier, suggesting that the magnitude gap traces cluster formation, which was also been found in N-body simulations of fossil group galaxies (D’Onghia et al., 2005; von Benda-Beckmann et al., 2008).
@@ 2.1 DES Y6 data
The data used throughout this analysis, both the source catalogs and the images, come from the DES Y6 co-addition (this includes publicly released data from DES Data Release 2 (Abbott et al., 2021)). The DESY6 images are deeper than the Y3 versions and provide a more uniform co-add with an improved background subtraction. Following the prescription described in Tanoglidis et al. (2021), the median surface brightness limits at 3 for the r- and i-band (the bands used in this analysis) are = and = mag arcsec-2, which is 0.5 mag arcsec-2 fainter than the DESY3 limits (Tanoglidis et al., 2021). Since this work focuses on the diffuse ICL, these improvements, which include the improved masking of faint galaxies located spatially near/within the ICL, are key. Such faint galaxies can unintentionally be treated as part of the ICL, thus, these improvements enhance our ability to accurately measure the contribution from the ICL by strengthening our ability to delineate between the diffuse light and total light within the cluster, which includes all cluster member galaxies and the ICL. We note that the improvements in the faint galaxy identification result from our use of the DES Y6 co-addition catalogues, which create a deeper catalogue by combining the full six years of r-,i-, and z-band photometry.
@@ 2.3 Measuring the ICL
We analyse only clusters in the redshift range 0.2 0.6, which reduces the number of redMaPPer ( 20) clusters from 21092 to 15654. This redshift range was selected due to the completeness range of DES-redMaPPer (Rykoff et al., 2016).
@@
Using these BCGs, the DES database was queried to select images within a 0.15 deg 0.15 deg region centred on the BCG 222For the redshift range in this study 0.15 deg ranges from 1.7Mpc - 3.6Mpc. These individual images were combined (using mean pixel values) to create the final coadd image centred on the BCG. The images used throughout this analysis have not had a local sky background subtraction (measured either with Source Extractor (Bertin & Arnouts, 1996) or Swarp (Bertin, 2010)) applied. Instead, these images underwent a global background subtraction process (estimated across the whole Field-of-View, about 2.5 for single exposure images; Bernstein et al., 2017). This background subtraction is done because local background regions often include the ICL, which results in an underestimation of the light in the BCG+ICL outer profiles (e.g., von der Linden et al., 2007; Golden-Marx & Miller, 2018).
Following the process described in Zhang et al. (2019c), we remove clusters with BCGs located within 2000 DES pixels (526″) from bright foreground stars and large nearby galaxies. This was done to remove saturated stars and yield a better sky estimate. These objects were identified using the bad region mask defined in Drlica-Wagner et al. (2018). Removing these clusters reduces the number of available clusters in our sample to 7042.
Using the BCG as the centre, all objects covering a 0.2 deg 0.2 deg region surrounding the BCG in the DES database were identified. We acknowledge that miscentring exists in redMaPPer (see Zhang et al. (2019a)). However, this is not corrected for in the reduction process. As shown in the Appendix, when we look at high and low probability centrals, we find no evidence that it impacts the parameters of the SMHM relation as shown in Tables 5 and 6.
Elliptical masks with a semi-major axis of 3.5 (Kron, 1980), were placed around all galaxies above a selected masking limit) excluding the BCG, allowing for the measurement of a BCG+ICL profile. The masking limit is a -band magnitude limit, given in Table 1, based on the cluster’s photometric redshift, of 0.2, where is the characteristic luminosity of a cluster red galaxy luminosity function measurement (Zhang et al., 2019b). We note that coadd catalogues are generally complete above 23.7 magnitudes in the z-band, so this tighter limit is done out of an abundance of caution.
Table 1: Satellite Galaxy Masking Limits redshift z-band magnitude limit 0.2 0.3 20.7 0.3 0.4 21.4 0.4 0.5 21.9 0.5 0.6 22.3 -
@@ 3 The Observable Measurements for Constructing the SMHM Relation
3.1 BCG and ICL Spatial Definitions
As noted in Section 1, it is observationally difficult to disentangle the light in the BCG from the ICL. Although we focus on the combined BCG+ICL light profile, for the purposes of the SMHM relation, we use the fixed radial apertures introduced in Zhang et al. (2024) to separate the BCG and ICL into 3 regions, 0-30 kpc (the BCG), 30-80 kpc (the transition region), and beyond 80kpc (the ICL). This choice of radial ranges is in excellent agreement with the individual ICL studies of Abell 85 (Montes et al., 2021) and the Perseus cluster (Kluge et al., 2024).
To further verify our choice for the inner region of the BCG, we fit the inner portion of our light profile to a single Sérsic profile, allowing us to measure the effective radius, Reff, of the BCG. In Figure 1, we compare the effective radius to both the halo mass estimated using redMaPPer richness (described in Section 3.3) and to M14 (described in Section 3.4). Though BCG size scales with halo mass, we see no trend in the effective radius with either halo mass or M14 due to the high mass range of our sample. Moreover, our selection of 30 kpc encapsulates two times the effective radius for the vast majority of our sample, which suggests that our choice of apertures generally captures the light of the BCG and ICL, respectively.
3.2 Stellar Mass
@@
Using the selected aperture magnitudes, we use the EzGal (Mancone & Gonzalez, 2012) SED modelling software to estimate the stellar mass. For EzGal, we assume a passive spectral model, as DES photometry prevents us from statistically constraining additional star formation parameters such as burst times and formation epochs. For converting magnitude to stellar mass, we assume the same parameters as Golden-Marx & Miller (2019), Golden-Marx et al. (2022), and Golden-Marx et al. (2023); a Bruzual & Charlot (2003) stellar population synthesis model, a Salpeter (1955) Initial Mass Function (IMF), a formation redshift of , and a metallicity of 0.008 (66% Z⊙). These parameters were selected to minimize the statistic between the measured and EzGal modeled photometry. Following Golden-Marx et al. (2022) and Golden-Marx et al. (2023), we assume a subsolar metallicity for each BCG+ICL system since observations find that the metallicity of early type galaxies decreases radially outward (e.g., Loubser & Sánchez-Blázquez, 2012; McDermid et al., 2015; Oliva-Altamirano et al., 2015; Edwards et al., 2024) and light from the ICL has been characterized as having a subsolar metallicity (Montes & Trujillo, 2018). Additionally, our choice of a passive spectral model is supported by observations which show that both BCGs (e.g., McDonald et al., 2016) and the ICL (e.g., McCabe et al., 2023) lack recent star formation. We also note that our stellar mass estimate is independent of our choice of formation redshift and that changing the Salpeter (1955) IMF to a Chabrier (2003) IMF uniformly lowers our stellar mass estimates by 0.25 dex. Thus, the SMHM relation’s amplitude is the only SMHM parameter impacted.
@@ 3.4 Magnitude Gap (M14)
M14 is defined uniformly throughout this paper, regardless of our choice of BCG+ICL aperture stellar mass. For the DES-redMaPPer clusters we identify cluster members as those with . This is less restrictive than Golden-Marx & Miller (2019) and Golden-Marx et al. (2022) since we do not match to the C4 catalog (Miller et al., 2005). This choice optimizes the number of clusters in our sample and only reduces the sample from 7042 to 6903 clusters. When we did this analysis with a threshold of our results were consistent.
@@
In Figure 2, we plot the M14 values against the DES-redMaPPer halo masses (estimated using Equation 1). Using 6 halo mass bins, we measure the median values of M14 as a function of halo mass (in red). The median value is consistent (within less than 1) across the entire range in halo mass. We note that the upper limit on M14 does slightly decrease with halo mass; however, this is likely a result of small number statistics and hierarchical growth (i.e., the most massive clusters are later forming (e.g., Matthee et al., 2017) and still evolving.)